OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x

OpenStudy (anonymous):

x=4 right?

OpenStudy (anonymous):

we have : x+20/2=3x have : x+20=6x x-6x=-20 -5x=-20 x=4 And right !:)